# MoVA: Adapting Mixture of Vision Experts to Multimodal Context

### v1

@@ 3.3 Fine-grained Expert Fusion with MoV-Adapter
We propose the MoV-Adapter to facilitate fine-grained expert representation extraction and integration based on multimodal context. As shown in Fig. 2, the MoV-Adapter consists of $L$ adapter blocks and a text encoder. Each block contains an expert knowledge extractor, a dynamic gating network, and a transformer block. For the $i$-th block, the input feature is denoted as $\mathbf{X}^{i}\in\mathbb{R}^{C\times H\times W}$ and we take the CLIP base encoder feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ as the input feature $\mathbf{X}^{1}$ of the first block. We use $\mathbf{G}$ to indicate the indices of chosen $K$ experts. The expert feature set is $\{\mathcal{F}_{j}|j\in\mathbf{G}\}$. The final output feature of the MoV-Adapter is connected to the LLM text embedding space by an MLP layer.
3.3.1 Text Encoder.
We introduce a pre-trained BERT as the text encoder to extract language context information from the user’s instruction. We take the [CLS] token from the output of the text encoder as the text token $\mathbf{X}_{T}\in\mathbb{R}^{C_{T}}$. It is worth noting that all the adapter blocks share the same text token.
Figure 2: The architecture of MoV-Adapter.
3.3.2 Expert Knowledge Extractor.
We adopt $N$ cross-attention layers as the expert knowledge extractor to achieve efficient knowledge extraction. Note that only the expert features $\{\mathcal{F}_{j}|j\in\mathbf{G}\}$ and their corresponding cross-attention layers are involved in the extraction. For each selected expert feature $\mathcal{F}_{j}\in\mathbb{R}^{C_{j}\times H_{j}\times W_{j}}$, we first align its resolution to $\mathbf{X}^{i}$ with bilinear interpolation:
@@
3.3.3 Dynamic Gating Network.
We employ a dynamic gating network to contribute to a fine-grained knowledge integration process for the conditional representation $\{\mathbf{Y}_{j}^{i}|j\in\mathbf{G}\}$. It is implemented with the softmax over the logits of an MLP layer, processing multimodal representation to generate expert-wise soft weight $\mathbf{P}^{i}\in\mathbb{R}^{K}$ for the output of each cross-attention layer in the extractor. Specifically, the input to the gating network is the concatenated vector of a visual token $\mathbf{X}^{i}_{V}\in\mathbb{R}^{C}$ and the text token $\mathbf{X}_{T}\in\mathbb{R}^{C_{T}}$. We obtain $\mathbf{X}^{i}_{V}$ with a global average pooling operation to $\mathbf{X}^{i}$. Then we concatenate them to compute the gating weights and the expert-wise outputs by computing the weighted sum:
@@
3.3.4 Transformer Block.
The transformer block in the adapter block follows the vanilla design, consisting of a self-attention layer and an FFN layer. Taking the fused visual representation $\hat{\mathbf{X}}^{i}$, its output will serve as the input feature $\mathbf{X}^{i+1}$ for the next adapter block.
Figure 3: The training strategy of MoVA. We enhance the task-specific knowledge extraction capacity of the MoV-Adapter in the first stage. Then, we excite model multimodal capacities in the supervised finetuning stage. The expert routing ability is unlocked in the last stage.
@@ 3.4 Training Paradigm
As depicted in Fig. 3, the training process of MoVA consists of three stages: MoV-Adapter pretraining, supervised finetuning, and expert-routing LoRA training.
3.4.1 Pretraining.
To improve multimodal generalization, we first construct 15M visual instruction samples across diverse public datasets for different downstream tasks as the training data:
  • Image Caption: DataComp-1B [17]  [^1] Only 4M image-text pairs are randomly selected for the efficiency, ShareGPT4V-PT [10], and ALLaVA-4V [8].
  • Visual Grounding and Localization: Objects365 [65], RefCOCO [81], VisualGenome [31], PointQA [53], and Flickr30K [58].
  • Chart Understanding: MMC-Instruction [44], Chart2Text [26], DVQA [25], and SciGraphQA [38].
  • Text Recognition and Document Parsing: LLaVAR-PT [85] and 3M English document images from Common Crawl  [^2] https://commoncrawl.org.
  • Biomedical Image Understanding: LLaVA-Med [36].
Note that for each dataset, we add the annotations of coarse-grained expert routing via the method proposed in Sec. 3.2.2. During the pretraining phase, we only optimize the MoV-Adapter along with the base vision encoder while preserving the capabilities of the initial large language model. Meanwhile, we leverage the routing annotations to choose experts and ignore representations from irrelevant ones during training.
3.4.2 Supervised Finetuning.
We utilize high-quality visual instruction tuning data that build upon LLaVA-665K [45] for finetuning. Additionally, we integrate several visual question answering datasets across various domains, such as DocVQA [56], ChartQA [54], InfographicVQA [55], AI2D [28], ST-VQA [6], TextVQA [66], SynthDoG-en [29], Geometry3K [51], PGPS9K [82], Geo170K [18], VQA-RAD [34], and SLAKE [42].
We also encompass equivalent comprehensive captions [10, 8, 33, 24] generated by the advanced GPT4-V [2] for improved world knowledge. In the supervised fine-tuning stage, task-specific vision experts are frozen and we jointly optimize the base vision encoder, MoV-Adapter, and LLM. The objective of supervised fine-tuning is to align the enhanced visual representation and the embedding of LLM, boosting its visual instruction-following capabilities. The coarse-grained routing annotations are also directly used in this training phase.
3.4.3 Expert-routing LoRA Training.
We introduce the expert-routing LoRA layers into the LLM and only train these LoRA layers in the final stage. We use the same instruction tuning data as the second stage for routing task tuning.
@@ 4.1 Implementation Details
As mentioned in Sec. 3.4, our training pipeline consists of three stages. In the pretraining stage, we use the AdamW optimizer with an initial learning rate of $2$$\times$$10^{-4}$, a batch size of 1024, and train the model for 1 epoch. We jointly finetune the weights of the base vision encoder, MoV-Adapter, and LLM with a batch size of 128 and an initial learning rate of $2$$\times$$10^{-5}$ during supervised fine-tuning. In the last stage, only the LoRA layers are trained and we keep the same hyperparameter setting as the supervised fine-tuning phase. We use 3 transformer blocks ($L$$=$$3$) in the MoV-Adapter and its hidden dimension is 1024, which is consistent with the base vision encoder CLIP. The input resolution of the base vision encoder is set as 672$\times$672. Two residual blocks with an average pooling are employed in the MoV-Adapter to reduce the number of output image tokens from 2304 to 576. More details about vision experts are released in the Appendix.
For the proxy setting performed in Tab. 1, we follow the default setting of LLaVA-1.5 but incorporate several additional datasets, including DocVQA [56], ChartQA [54], RefCOCO referring segmentation data [81], LLaVA-Med [36], VQA-RAD [34], and SLAKE [42].
4.2 MLLM Benchmarks
@@
LLaVA-1.5 [45] Vicuna-7B 558K 665K 576 1511/316 64.3 58.3 58.7 85.9
LLaVA-1.5 [45] Vicuna-13B 558K 665K 576 1531/295 67.7 63.6 62.1 27.6 7.6 85.9
@@
LLaVA-NeXT [46] Hermes-Yi-34B 558K 760K 2880 1631/397 79.3 79.0 46.5 87.7
@@
We empirically analyze the multimodal capacity and generalization ability of MoVA on a wide range of challenging MLLM benchmarks in Tab. 3. Specifically, this comprehensive assessment is conducted on MME [16], MMBench [49], QBench [76], MathVista [50], MathVerse [83], and POPE [39].
Compared to other open-source MLLMs with similar model complexity, MoVA with Vicuna-7B achieves the best performance across 7 MLLM benchmarks while offering a more favorable balance between training efficiency and performance. For instance, MoVA-7B surpasses the recent state-of-the-art LLaVA-NeXT-7B [46] with a dynamic high resolution design, processing only 20% image tokens.
Furthermore, we adopt Hermes-Yi-34B [1] as the LLM to validate the scaling property of MoVA. As depicted in Tab. 3, the performance of MoVA-34B is on par with popular proprietary MLLMs (e.g., Gemini-Pro [67]) and outperforms Qwen-VL-Plus [5] on 5 MLLM benchmarks. For example, MoVA establishes new records on MMBench and MMBench-CN, even surpassing the GPT-4V [2] by a clear margin. These results suggest that the ensemble of vision experts with adaptive expert routing can serve as an effective dimension for MLLM model scaling.
@@ 4.3 Visual Question Answering
Table 4: Performance comparison on VQA benchmarks. We present the number of model parameters of each MLLM for a clear complexity comparison. * denotes the methods with zero-shot setting.
@@
The evaluation results on VQA benchmarks are presented in Tab. 4. In this section, we divide these benchmarks into general VQA benchmarks [19, 23, 52] and text-oriented VQA benchmarks [66, 54, 56, 28].
Thanks to the dynamic and efficient task-specific knowledge extraction, MoVA achieves state-of-the-art performances across diverse VQA benchmarks. For general VQA benchmarks, MoVA-7B outperforms InternVL-Chat [12] equipped with InternViT-6B on VQAv2 [19] and GQA by 4.2% and 1.9%, respectively. Besides, MoVA shows its proficiency in text recognition in various scenarios, including scene text, chart, document, and diagram. For instance, MoVA-7B catches up to the current state-of-the-art generalist CogAgent [20] with 18 billion parameters on these text-oriented benchmarks with smaller model size. The MoVA model with 38B parameters even surpasses the well-established specialist model PALI-X-55B [11] by clear margins. The outstanding performances on distinct VQA benchmarks demonstrate MoVA’s robust generalization capabilities across diverse domains.
@@ 4.4 Visual Grounding
We conduct experiments on Referring Expression Comprehension (REC) benchmarks [81] to evaluate the visual grounding ability of MoVA. The results are presented in Tab. 5. Compared with the previous leading generalist CogVLM with 17B parameters, MoVA-7B attains higher scores on 6 of 8 splits while reducing the model size by 40%. Besides, the performance of MoVA-7B is on par with the state-of-the-art specialist models that are elaborately designed for grounding tasks. For example, MoVA-7B achieves a score of 90.22% on RefCOCO+ val, which is 2.46% higher than the score of UNINEXT-H [77]. Our largest model MoVA-34B further pushes the performance bound of visual grounding on these benchmarks. These impressive results demonstrate MoVA’s remarkable visual grounding capacity.
@@ 4.5 Medical Visual Question Answering
This experiment is conducted on popular medical VQA benchmarks VQA-RAD and SLAKE. We directly leverage the medical VQA evaluation metric adopted by LLaVA-Med. Each sample of VQA-RAD and SLAKE is observed only once during the training process of MoVA and LLaVA-1.5. For a fair comparison, we compare MoVA with the LLaVA-Med variant that is finetuned with only 1 epoch on the benchmark. The performance of the LLaVA-Med specialist that is fully finetuned on downstream tasks is also reported. As presented in Tab. 7, MoVA-7B consistently yields higher scores than LLaVA-Med and LLaVA-1.5 on both medical VQA benchmarks, exhibiting its medical visual chat ability.
4.6 Image Segmentation
In this experiment, we aim to investigate if task-specific knowledge can improve MoVA on the segmentation task. Therefore, we introduce a simple design to extend MoVA to segmentation tasks. Unlike segmentation generalists [32] that adopt an additional pixel decoder with high-resolution images for high-quality mask generation, we just formulate the referring segmentation task as sequential polygon generation [74]. We finetune MoVA and the baseline with a SAM-Huge [30] backbone on the RefCOCO referring segmentation datasets. MoVA achieves 57.1% gIoU on the testA benchmark, which is 2.6% higher than the 54.5% of baseline. This result indicates that MoVA is capable of exploiting task-specific knowledge to solve segmentation tasks.
@@ 4.6 Image Segmentation
Design POPE GQA ChartQA DocVQA
MoVA 88.6 64.8 68.3 81.3
Random routing 86.8 63.1 60.4 71.6
w/o routing 86.4 63.4 62.5 73.7
w/o MoV-Adapter 87.3 62.7 65.2 77.1
Table 8: Experimental results of $K$ varying from 1 to 3.
$K$ MMB POPE GQA ChartQA
Dynamic 70.4 88.6 64.8 68.3
1 68.4 89.2 64.0 64.9
2 71.0 87.6 63.5 66.7
3 68.1 85.7 63.2 67.4
@@ 4.7 Ablation Study
As presented in Tab. 7, we perform an ablation to thoroughly delve into the effect of each component. First, we try to replace the context-aware routing with random routing. Without task-relevant vision experts, the performance drops by a large margin, especially on the text-oriented ChartQA and DocVQA benchmarks. Removing context-aware routing to leverage all vision experts also brings similar results. It proves that both these modifications introduce biased information from irrelevant vision experts due to the removal of context-aware routing. Then, we ablate the effectiveness of the MoV-Adapter by replacing it with simple linear layers. The removal of fine-grained expert feature fusion downgrades performance across all datasets. These results delineate that each component in MoVA can consistently yield significant gains.
4.7.2 Number of activated experts.
@@
In the context-aware routing phase, the number of activated experts $K$ is dynamic. We compare such a data-dependent design with other variations of constant $K$ in this experiment. As presented in Tab. 9, the overall performance of dynamic $K$ consistently outperforms other models with constant $K$. This reveals this dynamic implementation can fully exploit the task-specific knowledge of relevant experts while avoiding the incorporation of biased information.
4.7.3 Adapter Design.
@@
In this section, we conduct ablation studies on the design of the MoV-Adapter. The number of adapter blocks directly affects the model’s complexity and efficiency. As presented in Table 9, we compared the impact of using 2, 3, and 4 adapter blocks on the model’s performance. We observed that the baseline with 3 blocks can achieve better performance than other settings with 2 blocks or 4 blocks. Therefore, we empirically set $L$ to 3 by default. Then, we substituted our multimodal gating for uniform gating to investigate its effectiveness. Each of the experts is assigned the same soft weight in the uniform gating. We find uniform gating brings consistent performance drops in the test benchmarks. It indicates that the lack of the dynamic soft-weight technique harms the overall performance since it fails to perform precise knowledge extraction.
4.8 Visualization
Figure 4: The coarse-grained and fine-grained routing visualization of MoVA.
In Fig. 4, we present a few samples in three different scenarios: visual grounding, chart understanding, and biomedical multimodal processing. For each example, we provide the question, image, routing information for the expert models, and their corresponding weights. The coarse-grained routing effectively identifies relevant experts for these cases. The fine-grained routing also accurately assigns specific weights to each identified expert, ensuring precise expert allocation.

### v2

@@ 3.3 Fine-grained Expert Fusion with MoV-Adapter
We propose the MoV-Adapter to facilitate fine-grained expert representation extraction and integration based on multimodal context. As shown in Figure 2, the MoV-Adapter consists of $L$ adapter blocks and a text encoder. Each block contains an expert knowledge extractor, a dynamic gating network, and a transformer block. For the $i$-th block, the input feature is denoted as $\mathbf{X}^{i}\in\mathbb{R}^{C\times H\times W}$ and we take the CLIP base encoder feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ as the input feature $\mathbf{X}^{1}$ of the first block. We use $\mathbf{G}$ to indicate the indices of chosen $K$ experts. The expert feature set is $\{\mathcal{F}_{j}|j\in\mathbf{G}\}$. The final output feature of $L$ adapter blocks is $\mathbf{X}^{L+1}$. Additionally, we apply two residual blocks [54] with an average pooling to $\mathbf{X}^{L+1}$ to obtain a coarser image feature $\mathbf{X}^{L+1}_{out}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$, which is further connected to the LLM text embedding space by an MLP layer.
Text Encoder. We introduce a pre-trained BERT as the text encoder to extract language context information from the user’s instruction. We take the [CLS] token from the output of the text encoder as the text token $\mathbf{X}_{T}\in\mathbb{R}^{C_{T}}$. It is worth noting that all the adapter blocks share the same text token.
Expert Knowledge Extractor. We adopt $N$ cross-attention layers as the expert knowledge extractor to achieve efficient knowledge extraction. Note that only the expert features $\{\mathcal{F}_{j}|j\in\mathbf{G}\}$ and their corresponding cross-attention layers are involved in the extraction. For each selected expert feature $\mathcal{F}_{j}\in\mathbb{R}^{C_{j}\times H_{j}\times W_{j}}$, we first align its resolution to $\mathbf{X}^{i}$ with bilinear interpolation:
@@
Figure 2: MoV-Adapter architecture.
Dynamic Gating Network. We employ a dynamic gating network to contribute to a fine-grained knowledge integration process for the conditional representation $\{\mathbf{Y}_{j}^{i}|j\in\mathbf{G}\}$. It is implemented with the softmax over the logits of an MLP layer, processing multimodal representation to generate expert-wise soft weight $\mathbf{P}^{i}\in\mathbb{R}^{K}$ for the output of each cross-attention layer in the extractor. Specifically, the input to the gating network is the concatenated vector of a visual token $\mathbf{X}^{i}_{V}\in\mathbb{R}^{C}$ and the text token $\mathbf{X}_{T}\in\mathbb{R}^{C_{T}}$. We obtain $\mathbf{X}^{i}_{V}$ with a global average pooling operation to $\mathbf{X}^{i}$. Then we concatenate them to compute the gating weights and the expert-wise outputs by computing the weighted sum:
@@
Transformer Block. The transformer block in the adapter block follows the vanilla design, consisting of a self-attention layer and an FFN layer. Taking the fused visual representation $\hat{\mathbf{X}}^{i}$, its output will serve as the input feature $\mathbf{X}^{i+1}$ for the next adapter block.
Figure 3: The training strategy of MoVA. We enhance the task-specific knowledge extraction capacity in the first stage. Then, we excite model multimodal capacities in the last stage.
@@ 3.4 Training Paradigm
As shown in Figure 3, the training process of MoVA consists of pretraining and supervised finetuning.
Pretraining. To improve multimodal generalization, we first construct 15M visual instruction samples across diverse domains as the training data: (i) Image caption data that covers 4M randomly selected samples from DataComp-1B [55], ShareGPT4V-PT [51], and ALLaVA-4V [56]. (ii) Visual grounding and localization dataset that encompasses Objects365 [57], RefCOCO [21], VisualGenome [58], PointQA [59], and Flickr30K [60]. (iii) Chart understanding data that includes MMC-Instruction [61], Chart2Text [62], DVQA [63], and SciGraphQA [64]. (iv) Text recognition and document parsing data that covers LLaVAR-PT [65] and 3M English document images from Common Crawl  [^2] https://commoncrawl.org. (v) LLaVA-Med [66] for biomedical image understanding. During the pretraining phase, we only optimize the MoV-Adapter along with the base vision encoder while preserving the capabilities of the initial large language model. Meanwhile, we leverage the routing annotations generated via the method proposed in Section 2 to choose experts and ignore representations from irrelevant ones during training.
Supervised Finetuning. We utilize high-quality visual instruction tuning data that build upon LLaVA-665K [28] for finetuning. Additionally, we integrate several visual question answering datasets across various domains, such as DocVQA [17], ChartQA [18], InfographicVQA [67], AI2D [68], ST-VQA [69], TextVQA [70], SynthDoG-en [71], Geometry3K [72], PGPS9K [73], Geo170K [74], RefCOCO, LLaVA-Med, VQA-RAD [75], and SLAKE [22]. We also encompass equivalent comprehensive captions [51, 56, 76, 77] generated by the advanced GPT4-V [53] for improved world knowledge. Apart from the above instruction tuning data, we convert the selected 2K routing annotations to instructions and incorporate them into the training data. In the supervised fine-tuning stage, only task-specific vision experts are frozen and we jointly optimize other components. The objective of supervised fine-tuning is to align the visual representation and the embedding of LLM, boosting its visual instruction-following capabilities.
@@ 4.1 Implementation Details
As mentioned in Section 3.4, our training pipeline consists of two stages. In the pretraining stage, we use the AdamW optimizer with an initial learning rate of $2$$\times$$10^{-4}$, a batch size of 1024, and train the model for 1 epoch. We jointly finetune the weights of all components except additional vision experts with a batch size of 128 and an initial learning rate of $2$$\times$$10^{-5}$ during supervised finetuning. We use 3 transformer blocks ($L$$=$$3$) in the MoV-Adapter and its hidden dimension is 1024, which is consistent with the base vision encoder CLIP. The input resolution of the base vision encoder is 672$\times$672. Two residual blocks with an average pooling are employed in the MoV-Adapter to reduce the number of output image tokens from 2304 to 576. For the experiment performed in Table 1, we follow the default setting of LLaVA-1.5 but incorporate several additional datasets, including DocVQA [17], ChartQA [18], RefCOCO [21], LLaVA-Med [66], VQA-RAD [75], and SLAKE [22]. More details about vision experts, ablations, and analysis are available in Appendix A.1 and A.3.
@@
LLaVA-1.5 [28] Vicuna-7B 558K 665K 576 1511/316 64.3 58.3 58.7 14.3 85.9
LLaVA-1.5 [28] Vicuna-13B 558K 665K 576 1531/295 67.7 63.6 62.1 27.6 17.0 85.9
@@
LLaVA-NeXT [80] Hermes-Yi-34B 558K 760K 2880 1631/397 79.3 79.0 46.5 23.4 87.7
@@
4.2 MLLM Benchmarks
We empirically analyze the multimodal capacity and generalization ability of MoVA on a wide range of challenging MLLM benchmarks in Table 3. This comprehensive assessment is conducted on MME [81], MMBench [16], QBench [82], MathVista [83], MathVerse [84], and POPE [20]. Compared to other open-source MLLMs with similar model complexity, MoVA with Vicuna-7B achieves the best performance across 7 MLLM benchmarks while offering a more favorable balance between training efficiency and performance. For instance, MoVA-7B surpasses the recent state-of-the-art LLaVA-NeXT-7B [80] with a dynamic high resolution design, processing only 20% image tokens. Furthermore, we adopt Hermes-Yi-34B [85] as the LLM to validate the scaling property of MoVA. As depicted in Table 3, the performance of MoVA-34B is on par with popular proprietary MLLMs (e.g., Gemini-Pro [78]) and outperforms Qwen-VL-Plus [6] on 5 MLLM benchmarks. For example, MoVA establishes new records on MMBench and MMBench-CN, even surpassing the GPT-4V [53]. These results suggest that the ensemble of vision experts with adaptive expert routing can serve as an effective dimension for MLLM model scaling.
@@ 4.3 Visual Question Answering
Table 4: Performance comparison on VQA benchmarks. We present the number of model parameters of each MLLM for a clear complexity comparison. * denotes zero-shot evaluation.
@@
The evaluation results on VQA benchmarks are outlined in Table 4. We divide these benchmarks into general VQA benchmarks [88, 19, 89] and text-oriented VQA benchmarks [70, 18, 17, 68]. Thanks to the dynamic and efficient task-specific knowledge extraction, MoVA achieves state-of-the-art performances across diverse VQA benchmarks. For general VQA benchmarks, MoVA-7B outperforms SPHINX-2k [4] equipped with Vicuna-13B on VQAv2 [88] and GQA by 4.2% and 1.9%, respectively. Besides, MoVA shows its proficiency in text recognition in various scenarios, encompassing scene text, chart, document, and diagram. For instance, MoVA-7B catches up to the current state-of-the-art generalist CogAgent [86] with 18 billion parameters on these text-oriented benchmarks with smaller model size. The MoVA model with 38B parameters even surpasses the well-established specialist model PALI-X-55B [87] by clear margins. These outstanding performances demonstrate MoVA’s robust generalization capabilities across diverse domains.
@@ 4.4 Visual Grounding
We conduct experiments on Referring Expression Comprehension (REC) benchmarks [21] to evaluate the visual grounding ability of MoVA. The results are presented in Table5. The performance of MoVA-7B is on par with the state-of-the-art specialist models that are elaborately designed for grounding tasks. For example, MoVA-7B achieves a score of 90.22% on RefCOCO+ val, which is 2.46% higher than the score of UNINEXT-H [92]. Our largest model MoVA-34B further pushes the performance bound of visual grounding on these benchmarks. These impressive results demonstrate MoVA’s remarkable visual grounding capacity.
@@ 4.5 Medical Visual Question Answering
This experiment is conducted on popular medical VQA benchmarks VQA-RAD and SLAKE. We directly leverage the medical VQA evaluation metric adopted by LLaVA-Med. Each sample of VQA-RAD and SLAKE is observed only once during the training process of MoVA and LLaVA-1.5. For a fair comparison, we compare MoVA with the LLaVA-Med variant that is finetuned with only 1 epoch on the benchmark. The performance of the LLaVA-Med specialist that is fully finetuned on downstream tasks is also reported. As presented in Table 8, MoVA-7B consistently yields higher scores than LLaVA-Med and LLaVA-1.5, exhibiting its medical visual chat ability.
@@
Design GQA ChartQA DocVQA
MoVA 64.8 68.3 81.3
Random routing 63.1 60.4 71.6
w/o routing 63.4 62.5 73.7
w/o MoV-Adapter 62.7 65.2 77.1
Table 8: Results of $K$ varying from 1 to 3.
$K$ GQA ChartQA
Dynamic 64.8 68.3
1 64.0 64.9
2 63.5 66.7
3 63.2 67.4
Table 9: Comparisons of expert routing criteria.
Design #Models POPE GQA ChartQA
Separate 4 88.6 64.8 68.3
Combination 14 88.9 64.6 68.7
Table 10: Open-world expert routing results.
Design #Samples Accuracy
MoVA 2K 92.5%
MoVA 5K 12.5%
w/o Augmentation 2K 0%
MLP classifier 2K 0%
@@ 4.6 Ablation Study
Component-wise analysis. As presented in Table 8, we perform an ablation to thoroughly delve into the effect of each component. First, we try to replace the context-aware routing with random routing. Without task-relevant vision experts, the performance drops by a large margin, especially on the text-oriented VQA benchmarks. Removing context-aware routing to leverage all vision experts also leads to similar results. It proves that both these modifications introduce biased information from irrelevant vision experts due to the removal of context-aware routing. Then, we ablate the effectiveness of the MoV-Adapter by replacing it with simple linear layers. The removal of fine-grained expert feature fusion downgrades performance across all datasets. These results delineate that each component in MoVA can consistently yield significant gains.
@@
Number of activated experts. In the context-aware routing phase, the number of activated experts $K$ is dynamic. We compare such a data-dependent design with other variations of constant $K$ in this experiment. As presented in Table 8, the overall performance of dynamic $K$ consistently outperforms other models with constant $K$. This reveals this dynamic implementation can fully exploit the task-specific knowledge of relevant experts while avoiding the incorporation of biased information.
Criteria for choosing better experts. To reduce the costs, our method only adopts ($N\choose 1$ models with $N$ various encoders to identify the better vision expert separately. However, we do not explicitly consider the combination of the chosen vision experts. In this experiment, we compare our method with another strategy that considers vision encoders combination and enumerates $\sum_{i=1}^{3}{N\choose i}$ models for routing data construction. Specifically, we set $N=4$ and employ a smaller model pool of DINOv2, Co-DETR, Pix2Struct, and Deplot to reduce training costs. As shown in Table 11, our method achieves comparable performance while requiring much less models for data construction.
Expert routing in open scenarios. We develop 105 human-verified testing samples that should be answered using novel experts for the expert routing task. These novel experts encompass 7 vision models [93, 71, 94, 91, 95, 54, 96] on various computer vision tasks and each expert corresponds to 15 evaluation samples. We manually check the correctness of the expert routing result. As presented in Table 11, a lightweight network, such as a MLP classifier fails to generalize to this open-world setting. Besides, increasing the routing training samples and removing data augmentations also lead to severe performance degradation. The results demonstrate our coarse-grained context-aware routing preserves the generalization ability for expert routing in open scenarios.
@@
Adapter Design. In this section, we conduct ablation studies on the design of the MoV-Adapter. As presented in Table 11, we compared the impact of using 2, 3, and 4 adapter blocks on the model’s performance. We observed that the baseline with 3 blocks can achieve better performance than other settings. Then, we substituted our multimodal gating for uniform gating to investigate its effectiveness. Each of the experts is assigned the same soft weight in the uniform gating. We find uniform gating brings consistent performance drops in the test benchmarks. It indicates that the lack of the dynamic soft-weight harms the overall performance since it fails to perform precise knowledge extraction.
Inference Analysis. As illustrated in Figure 1, MoVA consists of two stages: coarse-grained context-ware expert routing and fine-grained expert fusion with MoV-Adapter. This two-stage inference pipeline can be further broken down into five steps: (i) Data preprocessing. We first process the input image with image processors and convert the input text into a token sequence with the LLM tokenizer. (ii) Base encoder forward. We extract the base image feature using the base CLIP encoder. Note that we only run the base encoder once since its output feature can be preserved and reused in the fourth step. (iii) LLM routing generation. We compress the base image features into 144 image tokens. The LLM generates a concise routing answer based on the compressed image feature and routing instruction. Vision experts and MoV-Adapter forward. (iv) According to the multimodal context and routing results generated in the previous step, we fuse vision features of the base encoder and activated experts in a coarse-to-fine manner. (v) LLM response generation. The LLM generates the final response given the fused vision features and user instructions. To investigate the inference efficiency of each step, we randomly select 200 images from the COCO val2017 dataset and adopt the common image caption instruction: Describe this image. The temperature for generation is 0. The latency is measured using bfloat16 and flash-attention 2 on an A100 80G GPU. We present the average inference latency of each step and show the average sequence length of the routing output and final response. The inference latencies for each step are 0.19s, 0.05s, 0.14s, 0.07s, and 10.24s, respectively. The average length of the routing output is 3.24 tokens, while the average length of the final response is 405.06 tokens. Compared to the LLM response generation (Step 5), the LLM expert routing (Step 3) generates much fewer output tokens and its latency is negligible (0.14s v.s. 10.24s). Therefore, our method does not bring significant inference costs.